OpenStudy (anonymous):

Please help! How do I find the derivative of: the integral from 1 to sqrt(x) inverse tan(x) dx

OpenStudy (hunus):

\[d/dx\int\limits_{1}^\sqrt{x}\tan^{-1}(x)dx\] That it?

OpenStudy (anonymous):

Yep!

OpenStudy (anonymous):

it's inverse tan of x

OpenStudy (anonymous):

part of the fundamental theorem of calculus, the derivative of an an integral is the function that's being integrated

OpenStudy (hunus):

From the fundamental theorem of calculus we know that \[g(x)=\int\limits_{a=x}^{t=x}f(t)dt\] where a is a fixed a number and the upper limit of the integral varies \[g'(x)=f(x)\] Or the original functon f(t) evaluated at x For example \[g(x)=\int\limits_{1}^{x}\sqrt{1-t^2}dt\] \[g'(x)=\sqrt{1-x^2}\] When you take the derivative of an integral and the upper limit isn't just x (for example 2x, 3x, x^2) you must use the chain rule. We use the chain rule because we use substitution. \[g(x)=\int\limits_{1}^{x^2}\sin(t)dt\] We substitute x^2 = u \[g'(x)=d/dx \int\limits_{1}^{u}\sin(t)dt\] \[g'(x)=(d/du \int\limits_{1}^{u}\sin(t)dt) du/dx\] According to the fundamental theorem of calculus \[d/du\int\limits_{1}^{u}\sin{t}dt=\sin(u)\] This multiplied by du/dx gives \[du/dx(\sin(u))=2xsin(x^2)\] Now back to the original problem, using these methods \[d/dx\int\limits_{1}^{x^2}\tan^{-1}(t)dt\] \[2xtan^{-1}(x^2)\]

OpenStudy (hunus):

-.- I put the upper limit as x^2

OpenStudy (anonymous):

Thank you this is very helpful! However, why is the upper bound equal to x^(2)

OpenStudy (hunus):

Mistake on my part

OpenStudy (anonymous):

okay. thanks again!